Steven Parker owns and operates Steven's Septic Service and Legal Advice. Steven's two revenue generating (production) operations are supported by two service departments: Clerical and Janitorial. Costs in the service departments are allocated in the following order using the designated allocation bases:

Clerical:

Variable cost: expected number of work orders processed.
Fixed cost: long-run average number of work orders processed.

Janitorial:

Variable cost: labor hours.
Fixed cost: square footage of space occupied.
Average and expected activity levels for next month (June) are as follows:
  

No. of Work Orders
Expected Average Labor Hours Square Footage
Septic Service 45 80 710 1,950
Legal advice 30 25 990 2,350
Clerical 20 20 550 1,750
Janitorial 5 15 350 1,150


Expected costs in the service departments for June are as follows:

Clerical Janitorial
Variable costs $ 14,000 $ 5,700
Fixed costs $ 9,900 $ 950


Under the step method of allocation, how much Clerical Service cost should be allocated to the Septic Service operation for June? (Assume Clerical costs are allocated before Janitorial costs and round all calculations to the nearest whole dollar.)

rev: 08_19_2019_QC_CS-174994

  • $14,988.

  • $14,475.

  • $15,145.

  • $18,795.
Answer #1

Total expected number of work orders processed other than Clerical Service= 45+30+5= 80

Clerical service variable cost that should be allocated to Septic Service operation for June= $14000*45/80= $7875

Total average number of work orders processed other than Clerical Service= 80+25+15= 120

Clerical service fixed cost that should be allocated to Septic Service operation for June= $9900*80/120= $6600

Total Clerical Service cost should be allocated to the Septic Service operation for June= $7875+6600= $14475

So, the answer is option B) $14475